OpenStudy (anonymous):

Solve 3x2 − 24x = −12. 4 ± 2square root 2 4 ± 2square root 3 4 ± 2square root 5 4 ± 2square root 7

OpenStudy (nurali):

x² - 4x = - 4 + (- 4)² x² - 4x = - 4 + 16 (x - 4)² = 12 x - 4 = ± √12 OR ± 3.464101615 x = 4 ± √12 OR x = 7.464101615, 0.535898385 x = 4 ± 2√3 OR x = 7.464101615, 0.535898385

OpenStudy (anonymous):

can you do another one for me?

OpenStudy (nurali):

i will try.

OpenStudy (anonymous):

Solve x2 + 8x = 24.

OpenStudy (anonymous):

okay thank you!

OpenStudy (nurali):

Anytime.